Hash the values 94, 11, 39, 20, 16, 5, 12, 44, 13, 88 and 23 into a hash table with 11 slots. Handle collisions by linear probing. Use the hash function: h(i) (3 * i + 5) % 11 Type in each key on the same line as the index of the slot where it is eventually inserted. YOUR ANSWER TO QUESTION A: 0: 1: 2: 3: 5: 7: 9: 10:
Answer #1

hinear leicbing No. Date Tuseut 94 hi)-gy x3+5.11-1 Tnseut at wndek Twselt 1 hli) 1X3+5 11 = 5 tuekt ats 3) Tuseut 39 hli)- (2 No. Date Tweut y4 hti) 44x3+5) o 1-5 Next useut at 6 Besee aftes 5u 6 Alot aTuseut 13 hli) 13X3+5)/O Next feee dlot after oHash table 2 39 3 13 4% 23 11 6 8 7 88 8: 12 9: 16Nesl double Hashing his (3*its) /- LI h27-Ck 7) Tuseut 94 a4x3+5)7 11=L 16 embtt seut at 2 tnseut |1 (11x3+5)7-1-5 5 s empty5) wseut 16 (5+ 2xs 5t 315 buee dtet, dseut tuee St 4x5 5+515 54 645 6) Tuseut 5 (Sx3+5)/11 6ee 7-C97.7) S 9+5 711-3 9 nseut5+ 2x5) 4 not fuee (5+ 3x5)7 1= 1not 6uee 5+ yys 11 3 +5x5)-= 7 5+ 6ys)7 11= 2 at Buee nat buee Broe,nscut at 2 9 Tuseut 13 (Hasn table (double Htang O13 1:qu 2 44 3:S 4:39 723 8:12 9:16 l0: 20

if you like the answer please provide a thumbs up.